OpenStudy (anonymous):

Simplify the complex rational expression [ ( 6 / x-1 ) - ( 1 / x + 1) ] / [ ( x / x +1 ) + ( 1 / x + 1) ] into a standard rational expression.

OpenStudy (anonymous):

the denominator is 1

OpenStudy (anonymous):

Start by finding a common denominator for the numerator [6(x+1)/(x+1)(x-1)-(x+1)/(x+1)(x-1)] 6x+1-(x-1)/(x+1)(x-1) 5x+2/(x+1)(x-1) the denominator just = 1 5x+2/(x^2-1)

OpenStudy (anonymous):

\[\large \qquad \qquad \frac{\frac{6}{x-1} - \frac{1}{x+1}}{\frac{x}{x+1} + \frac{1}{x+1}}\]\[\large =\frac{\frac{6}{x-1} - \frac{1}{x+1}}{\frac{x}{x+1} + \frac{1}{x+1}} \cdot \frac{(x+1)(x-1)}{(x+1)(x-1)}\]\[\large =\frac{6(x+1) - (x-1)}{x(x-1) + (x-1)}\]\[\large = \frac{6x + 6 -x + 1}{x^2 - x + x - 1}\]\[\large =\frac{5x + 7}{x^2 - 1}\]

OpenStudy (anonymous):

Taco, you forgot to distribute your 6

OpenStudy (anonymous):

so It kept saying the dominator is incorrect.. :/

OpenStudy (anonymous):

Hrm.. You got the same thing?

OpenStudy (anonymous):

Yeah.. thats why it's saying the bottom one isn't right.

OpenStudy (anonymous):

That's why I'm confused..

OpenStudy (anonymous):

You wrote the problem incorrectly when you gave it to us.

OpenStudy (anonymous):

It should have been x-1 in the left had denominator at the bottom.

OpenStudy (anonymous):

Omg, I didn't even see that!

OpenStudy (anonymous):

Thanks for telling me!

OpenStudy (anonymous):

I wouldn't have known except that the screenshot showed the original problem ;)

OpenStudy (anonymous):

LOL.. sorry.. I'm an idiot.. let me see..let me try to solve this.

OpenStudy (anonymous):

Do you have skype, Polpak?

OpenStudy (anonymous):

Sure thing. Clerical errors always burn me too.

OpenStudy (anonymous):

GOT IT! Yess! Theehehee!!!!

OpenStudy (anonymous):

Thank you all!